# Logarithmic morphisms, tangential basepoints, and little disks

### v1

@@ 1.1. Overview
  • Various spaces and maps arising in quantum algebra and topology look like they should come from algebraic geometry, if only one could allow the maps to take values at infinity. A key example is the topological operad of little 2-disks. The space of $n$ little disks embedded in a bigger one is homotopy equivalent to the configuration space of $n$ distinct points in $\mathbb{C}$, which is an algebraic variety. However, the topological operation of gluing smaller disks into bigger ones is not expressible by a map of varieties. In this paper, building on work of Vaintrob [Vai21], we will explain how one can nevertheless realize the little disks operad algebro-geometrically—even over $\mathsf{Spec}(\mathbb{Z})$—as a “virtual logarithmic moduli space of genus zero curves with a tangential basepoint”, yielding a direct algebro-geometric proof of its formality, following a strategy of Beilinson [Bei00].
@@
The underlying conceptual issue is ultimately that tangential basepoints are not truly basepoints in the categorical sense, i.e. maps from the terminal object $*=\mathsf{Spec}(\mathbb{K})$ in the category of schemes over $\mathbb{K}$. The purpose of this paper is to explain how logarithmic algebraic geometry can be used to solve this problem. However, this requires an addition to the classical theory of log schemes of Fontaine–Illusie and Kato [Ill94, Kat89], which capitalizes on the following remarkable fact:
@@ 1.3. Tangential basepoints as virtual morphisms
Let $X$ be the log scheme associated to a strict normal crossing divisor $\underline{D}$ in a smooth variety $\underline{X}$ over a field $\mathbb{K}$. Let * denote the point $\mathsf{Spec}(\mathbb{K})$ equipped with the trivial log structure. Then virtual morphisms $*\to X$ of log schemes over $\mathbb{K}$ are in natural bijection with tangential basepoints for the pair $(\underline{X},\underline{D})$ in the sense of Deligne.
@@ 1.6. Virtual morphisms and little disks
Earlier proofs of this statement were given by Tamarkin [Tam03], Kontsevich [Kon99] with details by Lambrechts–Volić [LV14], Petersen [Pet14], Cirici–Horel [CH20], and Drummond-Cole–Horel [DCH21]. Our proof follows Beilinson’s suggestion [Bei00] that formality should be an immediate consequence of the fact that the little disks operad “comes from an operad in algebraic geometry”. Vaintrob [Vai21] previously managed to carry out that strategy for the framed little disks operad using his operad of log schemes and ordinary morphisms (Section 8.4); he then combined it with an extra argument about the behaviour of the cochain $\infty$-functor to deduce formality of the unframed little disks operad. Our proof, which needs virtual morphisms, is the first direct implementation of Beilinson’s suggestion.
@@ 1.7.3. Kähler differentials and smoothness
In the classical theory of schemes, there are many equivalent definitions of smoothness, but their natural analogues in log geometry do not all coincide. For instance, consider the standard log point $*_{\mathrm{log}}$ over a field $\mathbb{K}$. Its Kähler differentials $\Omega^{1}_{*_{\mathrm{log}}/\mathbb{K}}$ are free, and moreover, when $\mathbb{K}=\mathbb{C}$ its Kato–Nakayama space is a smooth manifold (the circle $\mathsf{S}^{1}$). However, $*_{\mathrm{log}}$ is not considered smooth over $\mathbb{K}$ (only “ideally smooth”, which is enough for our purposes). The reason is that in log geometry, “smoothness” is a condition on infinitesimal liftings, which cannot be solved in this case, because there are no ordinary morphisms $\mathsf{Spec}(\mathbb{K})\to*_{\mathrm{log}}$. In contrast, there are many virtual morphisms, which suggests that considering virtual infinitesimal liftings in the definition of smoothness may result in a theory in which the log point is smooth and the connection with Kähler differentials is more direct.
@@ 1.7.4. Fundamental groupoids
As illustrated by the example of parenthesized braids and little disks, the realization of tangential basepoints as virtual morphisms from a point allows one to give natural constructions of the fundamental groupoid with tangential basepoints in its various incarnations (de Rham, Betti, étale, etc.). This will be developed further in future work.
@@ 1.7.5. Riemann–Hilbert correspondence and “splittings”
Kato–Nakayama [KN99] and Ogus [Ogu03] have proven a Riemann–Hilbert correspondence for logarithmic flat connections on ideally smooth log schemes. This has recently been extended by Achinger [Ach23] to handle connections with regular singularities at infinity. A key ingredient in Achinger’s construction is the notion of a “splitting” of a log structure, which “behaves as if it were a section of the map from [a log scheme] $X^{\sharp}$ to its underlying scheme $\underline{X}^{\sharp}$ even though no such section exists, and in particular it allows one to turn log connections into classical connections by means of a ‘pull-back’ functor” (op. cit., p. 3). Virtual morphisms give a framework in which this intuition becomes precise: a splitting of an ideally smooth log scheme is literally equivalent to a section $\underline{X}^{\sharp}\to X^{\sharp}$ in the sense of virtual morphisms; see Section 2.4. The pullback on connections is then induced by the natural pullback on forms.
@@ 2.1. Log structures and log schemes
[…]
The log point $*_{\mathrm{log}}$ is the log scheme over $\mathbb{K}$ whose underlying scheme is $\mathsf{Spec}(\mathbb{K})$ and whose log structure is the pullback of the log structure of $\mathbb{A}^{1}_{\mathrm{log}}$ along the closed immersion $i\colon\{0\}\hookrightarrow\mathbb{A}^{1}$. Let $t$ denote the germ at 00 of the coordinate $z$ on $\mathbb{A}^{1}$. Then $\mathscr{M}_{*_{\mathrm{log}}}$ has global sections given by $\mathbb{K}^{\times}t^{\mathbb{N}}$, the product of $\mathbb{K}^{\times}$ with the free monoid generated by $t$, and $\alpha$ is given by “evaluation at $t=0$”:
@@
Equivalently, $*_{\mathrm{log}}=\mathsf{Spec}(t^{\mathbb{N}}\to\mathbb{K}[t]/(t))$. ∎
@@ 2.2. Ordinary and virtual morphisms
[…]
@@
[…]
@@
    Secondly, the pullback happens at the level of the group completions of the sheaves of monoids and produces formal quotients of monoid elements which are no longer associated to functions on the underlying scheme—this explains the terminology “virtual” (as in “virtual representation of a group”). For instance, for the log point $*_{\mathrm{log}}$ (Section 2.1), the monoid element $t$ corresponds to the function $\alpha(t)=0$ on the point, but its inverse $t^{-1}$ does not correspond to any function.
@@ 2.3. Categories of log schemes
In particular, if $\mathbb{K}$ is a ring and $S$ is the “point” $*:=\mathsf{Spec}(\mathbb{K})$ with the trivial log structure, we get the categories $\mathbf{LogSch}_{\mathbb{K}}$ and ${{}^{\mathbf{V}}}\mathbf{LogSch}_{\mathbb{K}}$ whose objects are log schemes over $\mathbb{K}$ in the usual sense.
@@ 2.4. Examples of virtual morphisms
[…]
Let $X$ be a log scheme over $\mathbb{K}$ and $*_{\mathrm{log}}$ be the log point (Section 2.1). A virtual morphism $f\colon X\to*_{\mathrm{log}}$ of log schemes over $\mathbb{K}$ is equivalent to the datum of a group homomorphism $t^{\mathbb{Z}}\to\mathscr{M}^{\mathrm{gp}}(X):=\mathsf{\Gamma}(X,\mathscr{M}^{\mathrm{gp}}_{X})$, i.e., an element $f^{*}(t)\in\mathscr{M}^{\mathrm{gp}}(X)$. Thus virtual morphisms $X\to*_{\mathrm{log}}$ of log schemes over $\mathbb{K}$ are in bijection with $\mathscr{M}^{\mathrm{gp}}(X)$. In contrast, ordinary morphisms $X\to*_{\mathrm{log}}$ are in bijection with $\mathscr{M}^{\mathrm{phan}}(X)\subset\mathscr{M}(X)$, the ideal of phantom global sections of $\mathscr{M}_{X}$, i.e. those sent to zero by $\alpha_{X}$. ∎
[…]
As a special case of the previous example, let $X=*$ denote the point $\mathsf{Spec}(\mathbb{K})$ equipped with the trivial log structure. There is a unique morphism $p\colon*_{\mathrm{log}}\to*$ of log schemes over $\mathbb{K}$, given by the identity map of $\mathsf{Spec}(\mathbb{K})$, with the map of monoids of global sections given by the inclusion $\mathbb{K}^{\times}\hookrightarrow\mathbb{K}^{\times}t^{\mathbb{N}}$. Meanwhile, a virtual morphism
@@
of log schemes over $\mathbb{K}$ is determined by an element $s^{*}(t)=\lambda\in\mathbb{K}^{\times}$. On global sections, $s^{*}\colon\mathbb{K}^{\times}t^{\mathbb{N}}\to\mathbb{K}^{\times}$ is the morphism of monoids given by evaluation at $t=\lambda$. Thus virtual morphisms $*\to*_{\mathrm{log}}$ of log schemes over $\mathbb{K}$ are in bijection with $\mathbb{K}^{\times}$, and are all sections of $p\colon*_{\mathrm{log}}\to*$. Note, however, that there are no ordinary morphisms from * to $*_{\mathrm{log}}$. Indeed, since $\alpha(t)=0$, an ordinary morphism would have to send $t$ to a phantom in $\mathbb{K}^{\times}$, of which there are none. ∎
[…]
A virtual automorphism $f$ of the log point $*_{\mathrm{log}}$ over $\mathbb{K}$ is equivalent to the datum of
@@
of log schemes over $\mathbb{K}$ is thus determined by an element $r^{*}(t)=\mu z^{j}$ with $\mu\in\mathbb{K}^{\times}$ and $j\in\mathbb{Z}$. If $\mu=1$ and $j=1$, then $r$ is a retraction of the ordinary morphism $i\colon*_{\mathrm{log}}\hookrightarrow\mathbb{A}^{1}_{\mathrm{log}}$ defined by $i^{*}(z)=t$. Note, however, that there are no ordinary morphisms from $\mathbb{A}^{1}_{\mathrm{log}}$ to $*_{\mathrm{log}}$ because $\mathscr{M}_{\mathbb{A}^{1}_{\mathrm{log}}}$ has no phantoms. ∎
[…]
@@
[…]
Let $X$ be a log scheme and let $\underline{X}$ be the underlying scheme, viewed as a log scheme with trivial log structure. Then there is a canonical ordinary morphism $\pi\colon X\to\underline{X}$, given by the inclusion $\mathcal{O}_{X}^{\times}\hookrightarrow\mathscr{M}_{X}$ of sheaves of monoids. When the log structure of $X$ is nontrivial, this morphism does not admit any ordinary sections, i.e. ordinary morphisms $s\colon\underline{X}\to X$ such that $\pi\circ s=\mathrm{id}_{\underline{X}}$. However it may admit virtual sections: these are equivalent to monoid homomorphisms $\mathscr{M}_{X}\to\mathcal{O}_{X}^{\times}$ that split the inclusion. For integral monoids, such a morphism is equivalent to a section of the projection $\mathscr{M}_{X}\to\mathscr{M}_{X}/\mathcal{O}_{X}^{\times}$; the latter were called “splittings” of the log structure in [Ach23]. ∎
@@ 2.5.2. Morphisms as monomial maps
[…]
@@ 2.5.3. Examples
[…]

### v2

@@ 1.1. Overview
  • Various spaces and maps arising in quantum algebra and topology look like they should come from algebraic geometry, if only one could allow the maps to take values at infinity. A key example is the topological operad of little 2-disks. The space of $n$ little disks embedded in a bigger one is homotopy equivalent to the configuration space of $n$ distinct points in $\mathbb{C}$, which is an algebraic variety. However, the topological operation of gluing smaller disks into bigger ones is not expressible by a map of varieties. In this paper, building on work of Vaintrob [Vai21], we will explain how one can nevertheless realize the little disks operad algebro-geometrically—even over $\mathsf{Spec}(\mathbb{Z})$—as a “virtual logarithmic moduli space of genus zero curves with a tangential basepoint”, yielding a direct algebro-geometric proof of its formality, following a strategy of Beilinson [Bei00]. The question of the algebro-geometric nature of the little disks operad was also raised independently by Morava [Mor07, appendix].
@@
The underlying conceptual issue is ultimately that tangential basepoints are not truly basepoints in the categorical sense, i.e. maps from the terminal object $\mathsf{Spec}(\mathbb{K})$ in the category of schemes over $\mathbb{K}$. The purpose of this paper is to explain how logarithmic algebraic geometry can be used to solve this problem. However, this requires an addition to the classical theory of log schemes of Fontaine–Illusie and Kato [Ill94, Kat89], which capitalizes on the following remarkable fact:
@@ 1.3. Tangential basepoints as virtual morphisms
Let $X$ be the log scheme associated to a strict normal crossing divisor $\underline{D}$ in a smooth variety $\underline{X}$ over a field $\mathbb{K}$. Let * denote the point $\mathsf{Spec}(\mathbb{K})$ equipped with the trivial log structure. Then virtual morphisms $\to X$ of log schemes over $\mathbb{K}$ are in natural bijection with tangential basepoints for the pair $(\underline{X},\underline{D})$ in the sense of Deligne.
@@ 1.6. Virtual morphisms and little disks
Earlier proofs of this statement were given by Tamarkin [Tam03], Kontsevich [Kon99] with details by Lambrechts–Volić [LV14], Petersen [Pet14], Cirici–Horel [CH20], and Drummond-Cole–Horel [DCH21]. Our proof follows Beilinson’s suggestion [Bei00] that formality should be an immediate consequence of the fact that the little disks operad “comes from an operad in algebraic geometry”. Vaintrob [Vai21] (see also [Vai19]) previously managed to carry out that strategy for the framed little disks operad using his operad of log schemes and ordinary morphisms (Section 8.4); he then combined it with an extra argument about the behaviour of the cochain $\infty$-functor to deduce formality of the unframed little disks operad. Our proof, which needs virtual morphisms, is the first direct implementation of Beilinson’s suggestion.
@@ 1.7.3. Kähler differentials and smoothness
In the classical theory of schemes, there are many equivalent definitions of smoothness, but their natural analogues in log geometry do not all coincide. For instance, consider the standard log point $_{\mathrm{log}}$ over a field $\mathbb{K}$. Its Kähler differentials $\Omega^{1}_{*_{\mathrm{log}}/\mathbb{K}}$ are free, and moreover, when $\mathbb{K}=\mathbb{C}$ its Kato–Nakayama space is a smooth manifold (the circle $\mathsf{S}^{1}$). However, $_{\mathrm{log}}$ is not considered smooth over $\mathbb{K}$ (only “ideally smooth”, which is enough for our purposes). The reason is that in log geometry, “smoothness” is a condition on infinitesimal liftings, which cannot be solved in this case, because there are no ordinary morphisms $\mathsf{Spec}(\mathbb{K})\to*_{\mathrm{log}}$. In contrast, there are many virtual morphisms, which suggests that considering virtual infinitesimal liftings in the definition of smoothness may result in a theory in which the log point is smooth and the connection with Kähler differentials is more direct.
@@ 1.7.4. Fundamental groupoids
As illustrated by the example of parenthesized braids and little disks, the realization of tangential basepoints as virtual morphisms from a point allows one to give natural constructions of the fundamental groupoid with tangential basepoints in its various incarnations (de Rham, Betti, étale, etc.). This will be developed further in future work. Note that in the context of $p$-adic Hodge theory, Olsson [Ols11, Chapter 9] already used logarithmic geometry to show how tangential basepoints provide fiber functors for relevant Tannakian categories.
@@ 1.7.5. Riemann–Hilbert correspondence and “splittings”
Kato–Nakayama [KN99] and Ogus [Ogu03] have proven a Riemann–Hilbert correspondence for logarithmic flat connections on ideally smooth log schemes. This has recently been extended by Achinger [Ach25] to handle connections with regular singularities at infinity. A key ingredient in Achinger’s construction is the notion of a “splitting” of a log structure, which “behaves as if it were a section of the map from [a log scheme] $X^{\sharp}$ to its underlying scheme $\underline{X}^{\sharp}$ even though no such section exists, and in particular it allows one to turn log connections into classical connections by means of a ‘pull-back’ functor” (op. cit., p. 3). Virtual morphisms give a framework in which this intuition becomes precise: a splitting of an ideally smooth log scheme is literally equivalent to a section $\underline{X}^{\sharp}\to X^{\sharp}$ in the sense of virtual morphisms; see Section 2.4. The pullback on connections is then induced by the natural pullback on forms.
@@ 2.1. Log structures and log schemes
Example 2.5 (The log point).
The log point $_{\mathrm{log}}$ is the log scheme over $\mathbb{K}$ whose underlying scheme is $\mathsf{Spec}(\mathbb{K})$ and whose log structure is the pullback of the log structure of $\mathbb{A}^{1}_{\mathrm{log}}$ along the closed immersion $i\colon\{0\}\hookrightarrow\mathbb{A}^{1}$. Let $t$ denote the germ at 0 of the coordinate $z$ on $\mathbb{A}^{1}$. Then $\mathscr{M}_{*_{\mathrm{log}}}$ has global sections given by $\mathbb{K}^{\times}t^{\mathbb{N}}$, the product of $\mathbb{K}^{\times}$ with the free monoid generated by $t$, and $\alpha$ is given by “evaluation at t=0t=0”:
@@
Equivalently, $_{\mathrm{log}}=\mathsf{Spec}(t^{\mathbb{N}}\to\mathbb{K}[t]/(t))$. ∎
@@ 2.2. Ordinary and virtual morphisms
$\underline{\phi}^{-1}\mathscr{M}_{Y}$$\mathscr{M}_{X}$$\underline{\phi}^{-1}\mathcal{O}_{Y}$$\mathcal{O}_{X}$$\scriptstyle{\phi^{*}}$$\scriptstyle{\alpha_{Y}}$$\scriptstyle{\alpha_{X}}$$\scriptstyle{\underline{\phi}^{*}}$ (2)
@@
$\underline{\phi}^{-1}\mathcal{O}_{Y}^{\times}$$\mathcal{O}_{X}^{\times}$$\underline{\phi}^{-1}\mathscr{M}^{\mathrm{gp}}_{Y}$$\mathscr{M}^{\mathrm{gp}}_{X}$$\scriptstyle{\underline{\phi}^{*}}$$\scriptstyle{\phi^{*}}$ (3)
Remark 2.7.
@@
    Secondly, the pullback happens at the level of the group completions of the sheaves of monoids and produces formal quotients of monoid elements which are no longer associated to functions on the underlying scheme—this explains the terminology “virtual” (as in “virtual representation of a group”). For instance, for the log point $_{\mathrm{log}}$ (Section 2.1), the monoid element $t$ corresponds to the function $\alpha(t)=0$ on the point, but its inverse $t^{-1}$ does not correspond to any function.
@@ 2.3. Categories of log schemes
In particular, if $\mathbb{K}$ is a ring and $S$ is the “point” $\mathsf{Spec}(\mathbb{K})$ with the trivial log structure, we get the categories $\mathbf{LogSch}_{\mathbb{K}}$ and ${{}^{\mathbf{V}}}\mathbf{LogSch}_{\mathbb{K}}$ whose objects are log schemes over $\mathbb{K}$ in the usual sense.
@@ 2.4. Examples of virtual morphisms
Example 2.10 (Virtual morphisms to a log point).
Let $X$ be a log scheme over $\mathbb{K}$ and $_{\mathrm{log}}$ be the log point (Section 2.1). A virtual morphism $f\colon X\to*_{\mathrm{log}}$ of log schemes over $\mathbb{K}$ is equivalent to the datum of a group homomorphism $t^{\mathbb{Z}}\to\mathscr{M}^{\mathrm{gp}}(X):=\mathsf{\Gamma}(X,\mathscr{M}^{\mathrm{gp}}_{X})$, i.e., an element $f^{*}(t)\in\mathscr{M}^{\mathrm{gp}}(X)$. Thus virtual morphisms $X\to*_{\mathrm{log}}$ of log schemes over $\mathbb{K}$ are in bijection with $\mathscr{M}^{\mathrm{gp}}(X)$. In contrast, ordinary morphisms $X\to*_{\mathrm{log}}$ are in bijection with $\mathscr{M}^{\mathrm{phan}}(X)\subset\mathscr{M}(X)$, the ideal of phantom global sections of $\mathscr{M}_{X}$, i.e. those sent to zero by $\alpha_{X}$. ∎
Example 2.11 (Virtual points of a log point).
As a special case of the previous example, let X=X=* denote the point $\mathsf{Spec}(\mathbb{K})$ equipped with the trivial log structure. There is a unique morphism $p\colon*_{\mathrm{log}}\to*$ of log schemes over $\mathbb{K}$, given by the identity map of $\mathsf{Spec}(\mathbb{K})$, with the map of monoids of global sections given by the inclusion $\mathbb{K}^{\times}\hookrightarrow\mathbb{K}^{\times}t^{\mathbb{N}}$. Meanwhile, a virtual morphism
@@
of log schemes over $\mathbb{K}$ is determined by an element $s^{*}(t)=\lambda\in\mathbb{K}^{\times}$. On global sections, $s^{*}\colon\mathbb{K}^{\times}t^{\mathbb{N}}\to\mathbb{K}^{\times}$ is the morphism of monoids given by evaluation at $t=\lambda$. Thus virtual morphisms $\to*_{\mathrm{log}}$ of log schemes over $\mathbb{K}$ are in bijection with $\mathbb{K}^{\times}$, and are all sections of $p\colon*_{\mathrm{log}}\to*$. Note, however, that there are no ordinary morphisms from * to $_{\mathrm{log}}$. Indeed, since $\alpha(t)=0$, an ordinary morphism would have to send $t$ to a phantom in $\mathbb{K}^{\times}$, of which there are none. ∎
Example 2.12 (Automorphisms of the log point).
A virtual automorphism $f$ of the log point $_{\mathrm{log}}$ over $\mathbb{K}$ is equivalent to the datum of
@@
of log schemes over $\mathbb{K}$ is thus determined by an element $r^{*}(t)=\mu z^{j}$ with $\mu\in\mathbb{K}^{\times}$ and $j\in\mathbb{Z}$. If $\mu=1$ and $j=1$, then $r$ is a retraction of the ordinary morphism $i\colon*_{\mathrm{log}}\hookrightarrow\mathbb{A}^{1}_{\mathrm{log}}$ defined by $i^{*}(z)=t$. Note, however, that there are no ordinary morphisms from $\mathbb{A}^{1}_{\mathrm{log}}$ to $_{\mathrm{log}}$ because $\mathscr{M}_{\mathbb{A}^{1}_{\mathrm{log}}}$ has no phantoms. ∎
Example 2.14 (Mapping a divisorial log structure to a log point).
@@
Example 2.15 (Splittings).
Let $X$ be a log scheme and let $\underline{X}$ be the underlying scheme, viewed as a log scheme with trivial log structure. Then there is a canonical ordinary morphism $\pi\colon X\to\underline{X}$, given by the inclusion $\mathcal{O}_{X}^{\times}\hookrightarrow\mathscr{M}_{X}$ of sheaves of monoids. When the log structure of $X$ is nontrivial, this morphism does not admit any ordinary sections, i.e. ordinary morphisms $s\colon\underline{X}\to X$ such that $\pi\circ s=\mathrm{id}_{\underline{X}}$. However it may admit virtual sections: these are equivalent to monoid homomorphisms $\mathscr{M}_{X}\to\mathcal{O}_{X}^{\times}$ that split the inclusion. For integral monoids, such a morphism is equivalent to a section of the projection $\mathscr{M}_{X}\to\mathscr{M}_{X}/\mathcal{O}_{X}^{\times}$; the latter were called “splittings” of the log structure in [Ach25]. ∎
@@ 2.5.2. Morphisms as monomial maps
Remark 2.18.
If $\underline{X}\setminus\underline{D}$ is a reduced scheme, then every map of $(\underline{X}\setminus\underline{D})$-schemes $\underline{E}^{\circ}|_{\underline{X}\setminus\underline{D}}\to\underline{E}^{\prime\circ}|_{\underline{X}\setminus\underline{D}}$ is automatically a monomial map, and therefore defines a virtual monomial map from $(\underline{X},\underline{D},\underline{E})$ to $(\underline{X},\underline{D},\underline{E}^{\prime})$. Indeed, if $R$ is a reduced ring, then the monomials $ft_{1}^{a_{1}}\cdots t_{n}^{a_{n}}$ with $f\in R^{\times}$ and $a_{i}\in\mathbb{Z}$ are the only invertible elements of the ring $R[t_{1}^{\pm 1},\ldots,t_{n}^{\pm 1}]$. This fails if $R$ is not reduced, e.g. $1+2t$ is its own inverse in $(\mathbb{Z}/4\mathbb{Z})[t^{\pm 1}]$. ∎
Proposition 2.19.
@@ 2.5.3. Examples
Example 2.23 (Virtual isomorphisms of Deligne–Faltings log structures).